# Private Data Leakage in Federated Human Activity Recognition for Wearable Healthcare Devices

### v1

@@ Abstract
Wearable wristband or watch can be utilized for health monitoring, such as determining the user’s activity status based on behavior and providing reasonable exercise recommendations. Obviously, the individual data perception and local computing capabilities of a single wearable device are limited, making it difficult to train a robust user behavior recognition model. Typically, joint modeling requires the collaboration of multiple wearable devices. An appropriate research approach is Federated Human Activity Recognition (HAR), which can train a global model without uploading users’ local exercise data. Nevertheless, recent studies indicate that federated learning still faces serious data security and privacy issues. To the best of our knowledge, there is no existing research on membership information leakage in Federated HAR. Therefore, our study aims to investigate the joint modeling process of multiple wearable devices for user behavior recognition, with a focus on analyzing the privacy leakage issues of wearable data. In our system, we consider a federated learning architecture consisting of $N$ wearable device users and a parameter server. The parameter server distributes the initial model to each user, who independently perceives their motion sensor data, conducts local model training, and uploads it to the server. The server aggregates these local models until convergence. In the federated learning architecture, the server may be curious and seek to obtain privacy information about relevant users from the model parameters. Hence, we consider membership inference attacks based on malicious servers, which exploit differences in model generalization across different client data. Through experimentation deployed on five publicly available HAR datasets, we demonstrate that the accuracy of malicious server membership inference reaches 92%. Our preliminary research confirms the significant privacy risks associated with joint training of multiple wearable devices, aiming to provide a novel research perspective in this field.
@@ II Motivation
II-A Human Activity Recognition (HAR)
HAR stands as a significant research domain within artificial intelligence and signal processing, aiming to utilize data collected from various sensors such as accelerometers and gyroscopes to identify and classify human activities [10]. HAR finds applications in diverse fields, including health monitoring, fall detection, and smart homes [11, 12, 13, 12]. Typically, HAR data is directly collected from sensors affixed to the human body, thereby increasing the potential risk of privacy exposure once malicious actors gain access to the data. Traditional HAR methods often rely on centralized data from different users or devices to train global models. However, this centralized approach raises concerns regarding privacy, security, and scalability [14].
@@ II-A Human Activity Recognition (HAR)
II-B Private Data Leakage In HAR
Figure 1: In the HARTH dataset, the walking data from subject 1 (left) and subject 2 (right) are represented, illustrating the signal graphs from the sensors placed on the back.
In many cases, federated learning addresses non-iid problems. However, there may exist varying privacy risks among the client devices involved in model training. In real-world application scenarios, HAR data collected from different individuals may exhibit differences due to individual habits, resulting in variations in the magnitude and posture of actions performed when executing the same activity. For instance, as depicted in Figure 1, the sensor data signals on the back in the walking activity data from subject 1 and subject 2 in the HARTH dataset exhibit significant differences despite representing the same activity. Consequently, due to the disparities in HAR data from different individuals, when training models using data from a single subject, the features learned by the trained model originate predominantly from that individual. As illustrated in Figure 2, when training the model using only subject 1’s data, the highest confidence in prediction vectors is consistently close to 1, and the distribution of data is similar. When subject 2’s data is input into the model trained with subject 1’s data, the highest confidence in the prediction vectors is only occasionally close to 1, with significant differences in the distribution of prediction vectors compared to those obtained from subject 1.
Figure 2: The model in the HARTH dataset is trained using data from subject 1. Specifically, 64 data points are randomly selected from both subject 1 and subject 2 datasets. These data points are then fed into the trained model to obtain prediction vectors, which are subsequently sorted in descending order. Only the top five confidence scores from each prediction vector are retained.
@@ III-A Federated HAR
Figure 3: The left part illustrates the process of HAR federated learning. The right part depicts the attacker saving the model $M_{j}$ uploaded by client $C_{j}$ during training, and using $M_{j}$ along with their own data to jointly train the attack model.
@@ IV-C HAR Model Performance
TABLE II: Average training and testing accuracy of five HAR datasets under various model conditions. The testing dataset comprises locally owned data from individual clients that were not involved in the training process.
@@
Figure 6: The accuracy and recall curves of the HAR70+ dataset under four distinct conditions of the 2$\times$ Conv model.
TABLE III: Accuracy and recall of the attack model trained under different conditions using the model uploaded by the target client for the first time on $D_{mix}$.
@@
TABLE IV: Accuracy and recall of the attack model trained on five HAR datasets under different conditions on $D_{mix}$.

### v2

@@ Abstract
Wearable data serves various health monitoring purposes, such as determining activity states based on user behavior and providing tailored exercise recommendations. However, the individual data perception and computational capabilities of wearable devices are limited, often necessitating the joint training of models across multiple devices. Federated Human Activity Recognition (HAR) presents a viable research avenue, allowing for global model training without the need to upload users’ local activity data. Nonetheless, recent studies have revealed significant privacy concerns persisting within federated learning frameworks. To address this gap, we focus on investigating privacy leakage issues within federated user behavior recognition modeling across multiple wearable devices. Our proposed system entails a federated learning architecture comprising $N$ wearable device users and a parameter server, which may exhibit curiosity in extracting sensitive user information from model parameters. Consequently, we consider a membership inference attack based on a malicious server, leveraging differences in model generalization across client data. Experimentation conducted on five publicly available HAR datasets demonstrates an accuracy rate of 92% for malicious server-based membership inference. Our study provides preliminary evidence of substantial privacy risks associated with federated training across multiple wearable devices, offering a novel research perspective within this domain.
@@ II Motivation
Human Activity Recognition (HAR): HAR stands as a significant research domain within artificial intelligence and signal processing, aiming to utilize data collected from various sensors such as accelerometers and gyroscopes to identify and classify human activities [10]. HAR finds applications in diverse fields, including health monitoring, fall detection, and smart homes [11, 12, 13, 12]. Typically, HAR data is directly collected from sensors affixed to the human body, thereby increasing the potential risk of privacy exposure once malicious actors gain access to the data. Traditional HAR methods often rely on centralized data from different users or devices to train global models. However, this centralized approach raises concerns regarding privacy, security, and scalability [14].
@@
Figure 1: The walking data of two subjects in the HARTH dataset.
Private Data Leakage In HAR: In many cases, federated learning addresses non-iid problems. However, there may exist varying privacy risks among the client devices involved in model training. In real-world application scenarios, HAR data collected from different individuals may exhibit differences due to individual habits, resulting in variations in the magnitude and posture of actions performed when executing the same activity. For instance, as depicted in Figure LABEL:fig:harth_c1_c2_walking, the sensor data signals on the back in the walking activity data from subject 1 and subject 2 in the HARTH dataset exhibit significant differences despite representing the same activity. Consequently, due to the disparities in HAR data from different individuals, when training models using data from a single subject, the features learned by the trained model originate predominantly from that individual. As illustrated in Figure 2, when training the model using only subject 1’s data, the highest confidence in prediction vectors is consistently close to 1, and the distribution of data is similar. When subject 2’s data is input into the model trained with subject 1’s data, the highest confidence in the prediction vectors is only occasionally close to 1, with significant differences in the distribution of prediction vectors compared to those obtained from subject 1.
Figure 2: The model trained from subject 1 in the HARTH dataset.
@@ III-A Federated HAR
Figure 3: The general framework of data leakage in Federated HAR.
@@ IV-C HAR Model Performance
TABLE II: Average training and testing accuracy under various model conditions.
@@
Figure 6: Attack performance of the HAR70+ dataset.
TABLE III: Attack performance on $D_{mix}$.
@@
TABLE IV: Attack performance on $D_{mix}$.